OpenStudy (anonymous):

solve the logarithmic equation: 10^(2x+5)=e

OpenStudy (anonymous):

\[\ln 2x+5=10\]

OpenStudy (anonymous):

\[ 10^{2x+5}=e \] to bring th exponent down, we take the logarithm to that base\[ \log_{10}10^{2x+5}=\log_{10}e\\ (2x+5)\log_{10}10=\log_{10}e\\ {\rm property:}\quad\log_aa=1\\ (2x+5)(1)=\log_{10}e\\ 2x=\log_{10}(e)-5\\ \vdots \]